The computer system a salesperson said I should get costs $3,525. He said that I could pay $1,500 down and make monthly payments of $100 per month for 2 years. My father said that he would sign for the loan since I am only sixteen. This computer has all the software I could ever want and has all the latest improvements.
For an additional $5 per month I can get a full service warranty for the two years that I make payments.
The printer is not included in the package. The salesperson said that trying to use the family printer might not work and would at the least be a nuisance to hook up each that that I wanted to use it. He suggested that I get a printer for $400.
What will be the total cost of the computer using the payment plan.

Answer by Theo(13342)

you don't know what the interest rate charge is, but you can find out using the online financial calculator at https://arachnoid.com/finance/index.html

input to this calculator to find the interest rate used would be:
pv = 3525 - 1500 = 2025
fv = 0
np = 24
pmt = -100
payment at end of time period

click on ir to get:
ir = 1.406336%
that's the interest rate per month.
to get the interest rate per year, multiply that by 12 to get:
1.406336% * 12 = 16.876032% interest rate per year.

here are the results from using that online calculator to find the interest rate per month.


if you pay for the printer with cash up front, the finance charge won't change.

if you pay for the printer on time, then your loan amount becomes 2025 + 400 = 2425.

with the same interest rate used for financing the computer, your payments per month will increase from 100 to 119.75 as shown below.

inputs to the calculator to find the increase in payments are:

pv = 2425
fv = 0
np = 24
ir = 1.406336%
payments at end of time period

click on pmt to get:
pmt = -119.75
that means you pay 119.75 at the end of each month instead of the original 100 at the end of each month for the computer alone.

here are the results of using the financial calculator for the loan that includes the additional 400 for the printer.



your original payment is 100 dollars a month, to which you would add the additional 5 dollars per month for the service warranty.

your revised payment is 119.75 dollars a month, to which you would add the additional 5 dollars per month for the service warranty on the computer.
if you buy insurance for the printer, you would add that to the monthly charge as well.

your total finance charge for the computer is 100 * 24 = 2400.
add your down payment of 1500 to get 3900.
add 5 * 24 = 120 for the insurance to get 4020.

your total finance charge for the computer and the printer is 119.75 * 25 = 2874.
add your down payment of 1500 for the computer to get 4374.
add 5 * 24 = 120 for the insurance on the computer to get 4494.
add whatever additional amount for the insurance on the printer, if you decide to buy insurance for that as well.

bottom line:

total payments of 4020 for the computer by itself plus insurance.

total payments of 44494 for the computer and the printer plus insurance on the computer. this does not include insurance for the printer.